# How Ambiguous Are the Rationales for Natural Language Reasoning? A Simple Approach to Handling Rationale Uncertainty

### v2

@@ Abstract
The quality of rationales is essential in the reasoning capabilities of language models. Rationales not only enhance reasoning performance in complex natural language tasks but also justify model decisions. However, obtaining impeccable rationales is often impossible. Our study aims to investigate how ambiguous rationales play in model performances of natural language reasoning. We first assess the ambiguity of rationales through the lens of entropy and uncertainty in model prior beliefs, exploring its impact on task performance. We then propose a simple way to guide models to choose between two different reasoning paths depending on the ambiguity of the rationales. Our empirical results demonstrate that this approach leads to robust performance, particularly in adversarial scenarios where rationale quality is inconsistent. [^1] The experiments were done at Classting AI Research, and the draft was mostly written at Oxford. The author is currently affiliated with the University of Oxford.
@@ 2.2 Ambiguous Rationales
The prior and posterior belief probabilities come from the model plausibility score as in Equation 1. We do not particularly measure model prior and posterior beliefs in the context of Bayesian settings but rather a probabilistic approach with a single model.
@@ 3 Experiments
In this section, we explain how we set up the experiments to observe the superiority of AURA which deals with aleatoric uncertainty in rationales for reasoning tasks. and what results we have observed.
@@ 3.1 Experimental Setup
We experiment with four question-answering datasets that require machine rationales in addition to one benchmark that includes human rationales: (a) CSQA (Talmor et al., 2019) a 5-choice dataset that requires commonsense knowledge extracted from ConceptNet. (a-1) ECQA (Aggarwal et al., 2021) explains each choice to questions from CSQA but with human rationales. (b) StrategyQA (Geva et al., 2021) is a dataset that requires an advanced level of reasoning strategies to produce binary (yes/no) answers. (c) OpenbookQA (Mihaylov et al., 2018) is a 4-choice dataset to test model ability on reasoning on more broad commonsense knowledge like open book testing. (d) QASC (Khot et al., 2020) is a complicated 8-choice question set crafted by a systematic composition of commonsense facts.
@@
We consider four rationalize-then-reason pipeline approaches similar to AURA as baselines. We also compare our approach with those with self-rationales or without any rationales. (1) Without rationales: fine-tune T5-base model without providing rationales. (2) Answer upon self-rationales using a medium-sized LM: make GPT-neox (20B) (Black et al., 2022) learn from a few examples to generate machine rationales and the following answers. (3) Answer upon self-rationales using a large-scale LM: same mechanism as (2) but use GPT3 (175B) for chain-of-thought reasoning Wei et al. (2022). (4) NILE (Kumar and Talukdar, 2020): a method of language inference over label-specific explanations. Following the standard setup, we fine-tune NILE with a T5 (3B) model, unlike other baselines that use a T5-base model. We present NILE only on the CSQA and StrategyQA datasets for fair comparison since other datasets do not provide equivalent setups for given golden rationales. (5) Train on prompted rationales with a standard training method: use GPT-neox but obtain answers not from the same model but from a separate reasoning model, T5-base in our experiment. (6) Dropout context is the same as (5) but randomly drops out from the input question contents while training the reasoning model (Wang et al., 2023). (7) same as (5) but trained with counterfactual regularization objective. We report the results presented by Wang et al. (2023).
@@ 3.2 Evaluation
We test AURA on In-Distribution (ID) and Out-of-Distribution (OoD) settings and different training ratios to investigate how AURA outperforms the existing methods.
@@
We test AURA on different training ratios to see its performance in low-resource settings. We observe that AURA is strong on limited resource settings on all four benchmarks as shown Figure 4. We observe AURA outperforms PINTO, the state-of-the-art rationalize-then-reason method, in both ID and OoD settings. Due to the nature of aleatoric uncertainty in rationales, the performance gains by AURA are pretty consistent over different training ratios.
@@ 4 Conclusion
Generating impeccable rationales is often impossible, and learning the patterns of rationales is almost improbable. We first explore how difficult it is for language models to learn rationales for reasoning tasks, both with human-annotated and machine-generated rationales. We then introduce a method of simply dealing with unclear machine-generated rationales, proposing a suitable two-system reasoning mechanism depending on the level of ambiguity. We empirically argue that AURA produces robust performance superiority against the adversarial quality of rationales and low-resource settings.

### v3

@@ Abstract
The quality of rationales is essential in the reasoning capabilities of language models. Rationales not only enhance reasoning performance in complex natural language tasks but also justify model decisions. However, obtaining impeccable rationales is often impossible. Our study aims to investigate how ambiguous rationales play in model performances of natural language reasoning. We first assess the ambiguity of rationales through the lens of entropy and uncertainty in model prior beliefs, exploring its impact on task performance. We then propose a simple way to guide models to choose between two different reasoning paths depending on the ambiguity of the rationales. Our empirical results demonstrate that this approach leads to robust performance, particularly in adversarial scenarios where rationale quality is inconsistent.
@@ A Simple Approach to Handling Rationale Uncertainty
**footnotetext: The work was mostly done at Classting AI Research. The author is currently affiliated with the University of Oxford.
@@ 2.2 Ambiguous Rationales
Model Prior and Posterior Beliefs.
LMs have shown that their parameters might contain a significant body of factual and commonsense knowledge as they get trained on a massive volume of text (Wei et al., 2022; Cohen et al., 2023). We call the internal knowledge that the models have had before being fine-tuned with the training set a prior belief of the model $\tilde{\theta}$. After the same model gets trained with a training set, we call their knowledge a posterior belief of the maximum likelihood model $\hat{\theta}_{\text{MLE}}$. The prior and posterior belief probabilities come from the model plausibility score as in Equation 1. We normalize the probability scores over the answer choices for each question. We do not particularly measure model prior and posterior beliefs in the context of Bayesian settings but rather a deterministic approach with a single model.
Rationale Ambiguity.
@@ 3 Experiments
In this section, we observe the superiority of AURA which deals with aleatoric uncertainty in rationales for reasoning tasks.
@@ 3.1 Experimental Setup
We experiment with four commonsense question-answering datasets that require machine rationales in addition to one benchmark that includes human rationales: (a) CSQA (Talmor et al., 2019) a 5-choice dataset that requires knowledge extracted from ConceptNet. (a-1) ECQA (Aggarwal et al., 2021) explains each choice to questions from CSQA but with human rationales. (b) StrategyQA (Geva et al., 2021) is a dataset that requires an advanced level of reasoning strategies to produce binary (yes/no) answers. (c) OpenbookQA (Mihaylov et al., 2018) is a 4-choice dataset to test model ability as a form of open book testing. (d) QASC (Khot et al., 2020) is a complicated 8-choice question set.
@@
We consider four rationalize-then-reason pipeline approaches similar to AURA as baselines. We also compare our approach with those with self-rationales or without any rationales. (1) Without rationales: fine-tune T5-base model without providing rationales. (2) Answer upon self-rationales using a medium-sized LM: make GPT-neox (20B) (Black et al., 2022) learn from a few examples to generate machine rationales and the following answers. (3) Answer upon self-rationales using a large-scale LM: same mechanism as (2) but use GPT3 (175B) Wei et al. (2022). (4) NILE (Kumar and Talukdar, 2020): a method of language inference on label-specific explanations. Following the standard setup, we fine-tune NILE with a T5 (3B) model, unlike other baselines that use a T5-base model. (5) Train on prompted rationales with a standard training method: use GPT-neox but obtain answers not from the same model but from a separate reasoning model, T5-base in our experiment. (6) Dropout context is the same as (5) but randomly drops out from the input question while training the reasoning model (Wang et al., 2023). (7) same as (5) but trained with counterfactual regularization objective. We report the results presented by Wang et al. (2023).
@@ 3.2 Evaluation
We test AURA on In-Distribution (ID) and Out-of-Distribution (OoD) settings and different training ratios to investigate the superiority of AURA.
@@
We test AURA on different training ratios to see its performance in low-resource settings. We observe that AURA is strong on limited resource settings on all four benchmarks as shown in Figure 4. We observe AURA outperforms PINTO, the state-of-the-art rationalize-then-reason method, in both ID and OoD settings. Due to the nature of aleatoric uncertainty in rationales, the performance gains by AURA are pretty consistent over different training ratios.
@@ 4 Conclusion
Generating impeccable rationales is often impossible, and learning the patterns of rationales is almost improbable. We first explore how ambiguous it is for language models to learn rationales for reasoning tasks, both with human-annotated and machine-generated rationales. We then introduce a method of simply dealing with unclear rationales, proposing a suitable two-system reasoning mechanism depending on the level of ambiguity. We empirically argue that AURA produces robust performance superiority against the adversarial quality of rationales and low-resource settings.
@@ 5 Limitations
Acknowledgements
This research was supported and funded by Classting Inc. The author thanks the members of the Classting AI Research, Jaesang Yoo, Yohaan Yoon, Cheoleui Hong, and Heeseok Jung, for their support for this work in various ways.